Find the critical value t* for the following situations.

a) 98 % confidence interval based on df=29.

b) 95 % confidence interval based on df=62.

a) What is the critical value of t for a 98 % confidence interval with df= 29?

b) What is the critical value of t for a 95% confidence interval with df=62?
Answer #1

Answer

(A) We have to find 98% confidence interval based on df = 29

Level of significance = 1 - (confidence level)/100 = 1 - (98/100) = 1-0.98 = 0.02

Using the degree of freedom 29 and level of significance 0.02 in the t distribution table

(look for df in the left most column and 0.02 in the top most row, select the intersection of both row and column)

we get

t critical = 2.462

(B) We have to find 95% confidence interval based on df = 62

Level of significance = 1 - (confidence level)/100 = 1 - (95/100) = 1-0.95 = 0.05

Using the degree of freedom 62 and level of significance 0.05 in the t distribution table

(look for df in the left most column and 0.05 in the top most row, select the intersection of both row and column)

we get

t critical = 1.999